# Towards Seamless Adaptation of Pre-trained Models for Visual Place Recognition

### 22 Feb 2024 v1

@@ 4.4 Ablation Study
The pre-trained DINOv2-GeM achieves decent results on Pitts30k, which has few dynamic objects on the place image. The performance is improved after full fine-tuning (i.e. Tuned-DINOv2-GeM), which indicates that even without dynamic interference, it is necessary to make the produced feature representation more suitable for VPR task through fine-tuning. However, Tuned-DINOv2-GeM performs worse than DINOv2-GeM on Tokyo24/7. It is because there is a generalization gap between Tokyo24/7 and Pitts30k-train, and full fine-tuning damages the excellent transferability of the pre-trained foundation model. Compared with Tuned-DINOv2-GeM, Global-Adaptation only tunes a few newly added adapters (with backbone frozen), which reduces the training consumption and always improves performance (retaining the generalization ability of the foundation model). Global-Adaptation achieves absolute R@1 improvement of 10.8% on Tokyo24/7. Besides, Local-Adaptation only adds a local adaptation module (with the proposed local loss for training) after the frozen DINOv2 to produce local features for re-ranking. It can also obviously improve performance, especially on Tokyo24/7 (20.6% absolute R@1 improvement), which shows day-night changes. The complete SelaVPR method combines the advantages of Glocal-Adaptation and Local-Adaptation, and achieves absolute R@1 improvement of 10% on Pitts30k and 26.3% on Tokyo24/7.
MSLS is a dataset with many dynamic objects (e.g. vehicles and pedestrians), on which the pre-trained DINOv2-GeM gets terrible results. By adding a local adaptation module after DINOv2 (Local-Adaptation) to get dense local features for re-ranking, we can get 22.1% absolute performance improvement for R@1, which is still not ideal. Using full fine-tuning can even bring significant performance improvement (34.6% for R@1), and the improvement yielded by Global-Adaptation is more obvious (42.3% for R@1). More importantly, the complete SelaVPR method achieves 2 × higher R@1.
@@ 5 Conclusions
In this paper, we introduced a novel hybrid adaptation method to seamlessly adapt pre-trained foundation models for the VPR task, which is composed of the global adaptation and local adaptation. The feature representation produced by the adapted foundation model is more focused on discriminative landmarks to differentiate places, thus bridging the gap between pre-training and VPR tasks. The local adaptation enables the model to output proper dense local features, which can be used for local matching re-ranking in two-stage VPR to greatly boost performance. The experimental results demonstrated that the proposed SelaVPR method outperforms previous SOTA methods on VPR benchmark datasets (with a large margin on Tokyo24/7 and Pitts30k). Meanwhile, since our approach eliminates the reliance on time-consuming spatial validation in re-ranking, it costs only 3% retrieval time of RANSAC-based two-stage methods. We believe that the proposed SelaVPR method provides a promising way to address the VPR task in real-world large-scale applications.
@@ Appendix A Comparisons of Tunable Parameters, Data Efficiency, and Training Time
Appendix B Performance of Different Global Features
This section compares the performance of using class token or GeM as global features in our SelaVPR method. As shown in Table 6, the performance of direct retrieval (without re-ranking) using class token as the global feature, i.e. SelaVPR-cls(global), is better than that of GeM. However, SelaVPR-GeM(re-rank) is better than SelaVPR-cls(re-rank) after re-ranking using local features. That is, the improvement of SelaVPR-cls after re-ranking is significantly lower than that of SelaVPR-GeM, and the performance of SelaVPR-cls after re-ranking is even decreased on MSLS-val. This shows that when we use local features in conjunction with global features, GeM is better compatible with local features, whereas class token is not. Therefore we finally choose GeM as the global feature in SelaVPR.
Appendix C Additional Ablation Experiments for the Global Adaptation
@@
Appendix D Comparisons on More Datasets
In this section, we compare the proposed method with other two-stage methods (with released code) on the Nordland and St. Lucia datasets. As shown in Table 9, our method surpasses all other methods even without re-ranking, i.e., SelaVPR(global). The SelaVPR outperforms other methods by a wide margin and achieves near-perfect results on St. Lucia (99.8% R@1 and 100% R@5).
Table 9: Comparison to two-stage methods on Nordland-test and St. Lucia.
@@ Appendix D Comparisons on More Datasets
Appendix E Qualitative Results of Local Matching
@@
Appendix F Additional Attention Visualization
Fig. 1 in the main paper shows that the feature representation of our method can precisely focus on image regions that are helpful for place recognition. Here, Fig. 7 demonstrates more challenging examples. Compared to the pre-trained model, which is disturbed by objects such as dynamic foreground, our method always pays attention to all discriminative regions (buildings and vegetation).
Figure 7: Additional attention map visualizations of the pre-trained model (DINOv2) and our model. We compute the mean in the channel dimension of the output feature map and display it using the heat map. Although in the last example our method is also affected by the bus, this is because it is closer and takes up a larger region of the image, while the buildings and vegetation are far away and occupy a smaller region. However, our method can still avoid missing any discriminative landmark (building and vegetation) to help retrieve correct results, while the pre-trained model cannot.
Appendix G Additional Qualitative Results
@@
Appendix H Dataset Details
@@
Pittsburgh (Torii et al., 2013). The Pittsburgh dataset is collected from Google Street View panoramas, and provides 24 images with different viewpoints at each place. The images in this dataset exist large viewpoint variations and moderate condition variations. We use Pitts30k in our experiments, which is a subset of Pitts250k (but more difficult than Pitts250k) and contains 10k database images each in the training, validation, and test sets.
@@
Appendix I Details of the Compared Method
@@
$R^{2}$ormer (Zhu et al., 2023) [^8] https://github.com/bytedance/R2Former. This work addresses both the retrieval and re-ranking in two-stage VPR using a novel transformer model. Its re-ranking module takes feature correlation, attention value, and coordinates into account, and does not require RANSAC-based geometric verification for re-ranking. $R^{2}$ormer is the SOTA two-stage VPR method. It ranked first in the MSLS challenge leaderboard before our method was proposed.
Appendix J The Snapshot of MSLS Leaderboard

### 18 Mar 2024 v2

@@ 4.4 Ablation Study
The pre-trained DINOv2-GeM achieves decent results on Pitts30k, which has few dynamic objects on the place image. The performance is improved after full fine-tuning (i.e. Tuned-DINOv2-GeM), indicating that fine-tuning is necessary to make the produced feature representation more suitable for the VPR task even without dynamic interference. However, Tuned-DINOv2-GeM performs worse than DINOv2-GeM on Tokyo24/7. It is because there is a generalization gap between Tokyo24/7 and Pitts30k-train, and full fine-tuning damages the excellent transferability of the pre-trained foundation model. Compared with Tuned-DINOv2-GeM, Global-Adaptation only tunes a few newly added adapters (with backbone frozen), which reduces the training consumption and always improves performance (retaining the generalization ability of the foundation model). Global-Adaptation achieves absolute R@1 improvement of 10.5% on Tokyo24/7. Besides, Local-Adaptation only adds a local adaptation module (with the proposed local loss for training) after the frozen DINOv2 to produce local features for re-ranking. It can also obviously improve performance, especially on Tokyo24/7 (20.3% absolute R@1 improvement), which shows day-night changes. The complete SelaVPR method combines the advantages of Glocal-Adaptation and Local-Adaptation, and achieves absolute R@1 improvement of 10.1% on Pitts30k and 26.0% on Tokyo24/7.
MSLS is a dataset with many dynamic objects (e.g. vehicles and pedestrians), on which the pre-trained DINOv2-GeM gets terrible results. By adding a local adaptation module after DINOv2 (Local-Adaptation) to get dense local features for re-ranking, we can get 22.5% absolute performance improvement for R@1, which is still not ideal. Using full fine-tuning can even bring significant performance improvement (35.0% for R@1), and the improvement yielded by Global-Adaptation is more obvious (42.7% for R@1). More importantly, the complete SelaVPR method achieves 2 × higher R@1.
@@ 5 Conclusions
In this paper, we introduced a novel hybrid adaptation method to seamlessly adapt pre-trained foundation models for the VPR task, which is composed of the global adaptation and local adaptation. The feature representation produced by the adapted foundation model is more focused on discriminative landmarks to differentiate places, thus bridging the gap between the pre-training and VPR tasks. The local adaptation enables the model to output proper dense local features, which can be used for local matching re-ranking in two-stage VPR to greatly boost performance. The experimental results demonstrated that the proposed SelaVPR method outperforms previous SOTA methods on VPR benchmark datasets (with a large margin on Tokyo24/7 and Pitts30k). Meanwhile, since our approach eliminates the reliance on time-consuming spatial validation in re-ranking, it costs only 3% retrieval time of RANSAC-based two-stage methods. We believe that the proposed SelaVPR method provides a promising way to address the VPR task in real-world large-scale applications.
@@ Appendix A Comparisons of Tunable Parameters, Data Efficiency, and Training Time
Appendix B Additional Ablation Experiments for the Global Adaptation
@@
Appendix C Performance of Different Global Features
This section compares the performance of using class token or GeM as global features in our SelaVPR method. As shown in Table 8, the performance of direct global retrieval (without re-ranking) using class token as the global feature, i.e. SelaVPR-cls(global), is better than that of GeM. However, SelaVPR-GeM(re-rank) is better than SelaVPR-cls(re-rank) after re-ranking using local features. That is, the improvement of SelaVPR-cls after re-ranking is significantly lower than that of SelaVPR-GeM, and the performance of SelaVPR-cls after re-ranking is even decreased on MSLS-val. This shows that when we use local features in conjunction with global features, GeM is better compatible with local features, whereas class token is not. Therefore we finally choose GeM as the global feature in SelaVPR.
Table 8: Performance of different global features without or with re-ranking.
Method Pitts30k-test MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10
SelaVPR-cls(global) 89.2 96.2 97.6 89.3 96.2 97.2
SelaVPR-GeM(global) 87.2 95.1 97.2 87.7 95.8 96.6
SelaVPR-cls(re-rank) 90.1 96.0 97.4 83.5 92.2 94.1
SelaVPR-GeM(re-rank) 91.4 96.5 97.8 90.8 96.4 97.2
Table 9: The results of using “local” features at different granularities for re-ranking.
Method Pitts30k-test MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10
global adaptation for global retrieval 87.3 94.6 96.6 87.4 95.9 96.9
coarse patch tokens for re-ranking 89.8 95.4 96.9 82.8 92.0 94.9
dense local features for re-ranking 91.4 96.5 97.8 90.8 96.4 97.2
Appendix D Performance of “Local” Features at Different Granularities
To illustrate the necessity of using dense local features in reranking, we provide a comparison of reranking using coarse 16$\times$16 patch tokens (directly produced by the global adapted backbone, and treated as coarse-grained “local” features) and our dense local features. The results are shown in Table 9. Although re-ranking with coarse patch features provides an improvement (but not as good as ours) on Pitts30k compared to only global adaptation for global retrieval (without re-ranking), it performs significantly worse than global retrieval on MSLS-val, where VPR methods are more susceptible to perceptual aliasing. That is, it will bring negative effects and damage the results of global retrieval. However, using dense local features (produced by local adaptation) for re-ranking can always provide significant improvements compared to global retrieval.
Appendix E Performance of Re-ranking Different Numbers of Candidates
Table 10: The results of re-ranking different numbers of candidates. Note that the model tested on Tokyo24/7 and Pitts30k-test is fine-tuned on MSLS-train and then further fine-tuned on Pitts30k-train (same as compared with the SOTA methods, but different from the ablation experiments).
Candidates Tokyo24/7 Pitts30k-test MSLS-val
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
Top-20 93.0 96.2 97.1 92.7 96.6 97.5 90.7 96.4 97.0
Top-50 93.7 96.5 96.8 92.8 96.8 97.6 90.8 96.4 97.0
Top-100 94.0 96.8 97.5 92.8 96.8 97.7 90.8 96.4 97.2
Top-200 94.0 96.8 97.5 92.8 96.9 97.8 90.7 96.5 97.3
We show the results of our SelaVPR for re-ranking different numbers of candidates in Table 10. The performance roughly reaches saturation when re-ranking top-100 candidates, which is our recommended setting for optimal recognition performance. Our approach also achieves excellent performance when only the top-20 candidates are re-ranked. This setting can reduce the re-ranking runtime by approximately 80%.
Appendix F Comparisons on More Datasets
This section provides comparisons with other methods on more datasets. We compare the proposed method with other two-stage methods on the Nordland and St. Lucia datasets. As shown in Table 11, our method surpasses other methods even without re-ranking, i.e., SelaVPR(global). The SelaVPR outperforms other methods by a wide margin and achieves near-perfect results on St. Lucia (99.8% R@1 and 100% R@5). We also provide the results of our SelaVPR on Pitts250k in Table 12. MixVPR achieves the previous best results on Pitts250k, and SelaVPR outperforms it.
Table 11: Comparison to two-stage methods on Nordland-test and St. Lucia. The model is the same as that used for MSLS in Table 2.
@@ Appendix F Comparisons on More Datasets
Table 12: The results of MixVPR and our method on Pitts250k. The model is the same as that used for Pitts30k in Table 2.
Method Pitts250k-test
R@1 R@5 R@10
MixVPR 94.3 98.2 98.9
SelaVPR 95.7 98.8 99.2
Appendix G Qualitative Results of Local Matching
@@
Appendix H Additional Attention Visualization
Fig. 1 in the main paper has shown that the feature representation of our method can precisely focus on image regions that are helpful for place recognition. Here, Fig. 7 demonstrates more challenging examples. Compared to the pre-trained model, which is disturbed by objects such as dynamic foreground, our method always pays attention to all discriminative regions (buildings and vegetation).
Figure 7: Additional attention map visualizations of the pre-trained model (DINOv2) and our model. We compute the mean in the channel dimension of the output feature map and display it using the heat map. In the last example, although our method is also affected by the bus, this is because it is closer and takes up a larger region of the image, while the buildings and vegetation are far away and occupy a smaller region. However, our method can still avoid missing any discriminative landmark (building and vegetation) to help retrieve correct results, while the pre-trained model cannot.
Appendix I Additional Qualitative Results
@@
Appendix J Dataset Details
@@
Pittsburgh (Torii et al., 2013). The Pittsburgh dataset is collected from Google Street View panoramas, and provides 24 images with different viewpoints at each place. The images in this dataset exist large viewpoint variations and moderate condition variations. The Pitts30k dataset is a subset of Pitts250k (but more difficult than Pitts250k) and contains 10k database images each in the training, validation, and test sets.
@@
Appendix K Details of the Compared Methods
@@
$R^{2}$ormer (Zhu et al., 2023) [^8] https://github.com/bytedance/R2Former. This work addresses both the retrieval and re-ranking in two-stage VPR using a novel transformer model. Its re-ranking module takes feature correlation, attention value, and coordinates into account, and does not require RANSAC-based geometric verification. $R^{2}$ormer is the SOTA two-stage VPR method. It ranked first in the MSLS challenge leaderboard before our method was proposed.
Appendix L The Snapshot of MSLS Leaderboard